It takes you 3 hours to drive to a concert 135 miles away. You drive 55 miles per hour on highways and 40 miles per hour the rest of the time. How much time did you spend driving at each speed?
Answer by robertb(5830)

x = time on h-ways, y = rest of the time.
Then x+y = 3 and 55x + 40y = 135. By direct substitution,
55x + 40(3-x) = 135,
55x+ 120-40x = 135,
15x = 135-120,
15x = 15,
x = 1, he drove 1 hr on h-ways, and 2 hours others.